OpenStudy (whitelightning98):

A data set consists of the following data points:(3, 5), (5, 8), (6, 13)The line of best fit has the equation = -3 + 2.5x. What does this equation predict for a value of x = 3?

OpenStudy (calculusxy):

is the question saying that the three points shown above all share the same slope-int formula: y = -3 + 2.5x ?

OpenStudy (whitelightning98):

i think so

OpenStudy (calculusxy):

well we have a coordinate that has the x-coordinate of 3 already (3,5) but if we use that formula we get: y = -3 + 2.5(3) -> y = -3 + 7.5 -> y = 4.5

OpenStudy (whitelightning98):

A. 7.5 B. 6 C. 4.5 D. 9 these are the multiple choice answers

OpenStudy (calculusxy):

okay so we have 4.5 as y.

OpenStudy (calculusxy):

recall the work that i did in the previous post. \(y = -3 + 2.5x\) \(y = -3 + 2.5(3)\) \(y = -3 + 7.5\) \(\large \color{green}{y = 4.5}\)

OpenStudy (whitelightning98):

so is the answer 4.5

OpenStudy (calculusxy):

yes

OpenStudy (calculusxy):

and you have the work as proof and my thought process :)

OpenStudy (whitelightning98):

thank you so much

OpenStudy (calculusxy):

you're welcome :)